OpenStudy (anonymous):

4x + 2(x − 2) = 4x + x − 12

OpenStudy (anonymous):

What is the colution for this equation?

OpenStudy (anonymous):

I think im putting too much thought into it

OpenStudy (anonymous):

x=-8

OpenStudy (anonymous):

How did you get that answer? Im trying to solve it too

OpenStudy (anonymous):

So i can know how to solve my other questions

OpenStudy (anonymous):

Step 1: Simplify both sides of the equation. 4x+2(x−2)=4x+x−12 Simplify: 6x−4=5x−12 Step 2: Subtract 5x from both sides. 6x−4−5x=5x−12−5x x−4=−12 Step 3: Add 4 to both sides. x−4+4=−12+4 x=−8

OpenStudy (anonymous):

Hope this was helpful :)

OpenStudy (anonymous):

This is Hard O_O

OpenStudy (anonymous):

I need like a algebra tutor or something

OpenStudy (anonymous):

I kinda understand a bit b etter now but im still stuck in space

OpenStudy (anonymous):

Believe me it wasn't easy getting through Algebra 1 myself lol

OpenStudy (anonymous):

I got an online tutor